Using the table, create a multiple linear regression model evaluating factor 1[x1] and factor 2[x2] against the response y.

| 26 1.0 1.5 175 160 16.3 4.0 2.0 100 0.5 3.0 1.0 10.5 | 1.0

Also, calculate the coefficient of determination and complete the regression ANOVA to determine if the model is valid.
Answer #1

Here I attach the R code and output

y=c(26,24,175,160,163,55,62,100,26,30,70,71)
x1=c(1,1,1.5,1.5,1.5,0.5,1.5,0.5,1,0.5,1,0.5)
x2=c(1,1,4,4,4,2,2,3,1.5,1.5,2.5,2.5)
fit=lm(y~x1+x2)
summary(fit)

anova(fit)

1586628066351_image.png

since the p value for both the variables are less than 0.05 then we can say that both x1 and x2 are significant

The regression model can be written as y=-49.635+18.355*x1+46.116*x2

the coefficient of determination R squared is 0.9771 which means that 97.71 percent of the total variation in the regression can be explained by the regression model.

The ANOVA table is given by

1586628365218_image.png